OpenStudy (anonymous):

If F(x) = f(x)g(x), where f and g have derivatives of all orders, show that F'' = f''g + 2f'g' + fg''

myininaya (myininaya):

you have a product so use the product rule \[(RS \cdot T)'=(RS)'\cdot (T)+(RS) \cdot (T)'\] R, S, and T are functions of x \[(R \cdot \S)'=(R)' \cdot (\S)+(R) \cdot (\S)'\]

myininaya (myininaya):

i couldn't get those things to look like S's

OpenStudy (zarkon):

F'=f'g+fg' F''=(f'g+fg')'=(f'g)'+(fg')' =(f''g+f'g')+(f'g'+fg'')=f''g+2f'g'+fg''